I dont even know where to start
Find the value of the polynomial x3 – 4x when x = –2.
A) –2
B) 0
C) 2
D) 14

Answer by edjones(8007)

x3 – 4x
Start by substituting -2 for x.
-2^3-4*-2=-8+8=0
Ed